OpenStudy (anonymous):

Simplify this rational expression 7m+14/5m+10 Show steps please! I need help!

OpenStudy (anonymous):

7m + 14 ------------- 5m + 10 7(m + 2) ------------- {factored 7 out of top and 5 out of bottom} 5(m + 2) 7 ---- {cancelled (m + 2) on top and bottom} 5

OpenStudy (anonymous):

thanks!!!!! im struggling with this!

OpenStudy (anonymous):

i have a similar problem and the answer came out to 8/4 do i simplify that or does that end the problem?

OpenStudy (anonymous):

the problem is 8z-24 ------ 4Z-12

OpenStudy (anonymous):

yes your answer is correct 8/4=2 you have to follow the same mathematical procedure

OpenStudy (anonymous):

THANKS!

OpenStudy (anonymous):

You are most welcomed

hero (hero):

\[\frac{7m + 14}{5m + 10} = \frac{7(m + 2)}{5(m + 2)} = \frac{7}{5} = \frac{5 + 2}{5} = \frac{5}{5} + \frac{2}{5} = 1 + \frac{2}{5}\]

hero (hero):

7/5 is improper fraction so reduce it to mixed fraction.